Find the missing amounts if cash is 2100,amount receivable- ? , Equipment-5000, bank loan payable - 2000, amounts payable- 750 begging capital-3000 drawing- 1200 revenue - 10000, expenses- 7000 , capital at the end - ?
1
Expert's answer
2021-09-09T17:14:13-0400

Solution:

Assets = Liabilities + Owners Equity

Total assets = Cash + amount receivable + Equipment

Total liabilities = Bank loan payable + amounts payable

Owners Equity (capital at the end) = Beginning capital + Net income – Drawings

Derive Net income:

Net income = Revenue – Expenses = 10,000 – 7000 = 3000

Owners Equity = 3,000 + 3,000 – 1,200 = 4,800

Capital at the end = 4,800


Total liabilities = 2,000 + 750 = 2,750

Total liabilities + Owners Equity = 2,750 + 4,800 = 7,550

Amount receivable = 7,550 – 2,100 – 5,000 = 450


Amount receivable = 450